A bus travels 225km in 5hours find the average speed in m/s
1
Expert's answer
2019-10-25T11:02:53-0400

The average speed of the bus:


"v=\frac{s}{t}"

"v=\frac{225000}{5\cdot 3600}=12.5\frac{m}{s}"